ZEEMAN TUNING RATES IN THE ν_{3} BAND OF NO_{2}

\underline{C. R. MAHON}, Space Physics Research Institute, 572 Hyannis Dr., Sunnyvale, CA 94087; C. CHACKERIAN, JR., L. P. GIVER, NASA Ames Research Center, Moffett Field, CA 94035; T. A. BLAKE, Battelle/PNL, P. O. Box 999, MS K3-58, Richland, WA 99352.